# Testing the Molecular Cloud Paradigm for Ultra-High-Energy Gamma Ray Emission from the Direction of SNR G106.3+2.7

### v1

@@ Abstract
Aims. We aim to characterize the very-high-energy (VHE, 0.1–100 TeV) gamma ray emission from SNR G106.3+2.7 by determining the morphology and spectral energy distribution of the region. This is accomplished using 2565 days of data and improved reconstruction algorithms from the HAWC Observatory. We also explore possible gamma ray production mechanisms for different energy ranges.
@@ 1 Introduction
SNR G106.3+2.7 is a PeVatron candidate that has been under study for many years now. It was first detected in 1990 by Joncas & Higgs (1990) at 408 MHz with the Dominion Radio Astrophysical Observatory (DRAO). The head of the SNR is home to the pulsar (PSR) J2229+6114 and its boomerang-like pulsar wind nebula (PWN) (Kothes et al., 2001), known as the Boomerang region. An extended shell then reaches into the tail region of the SNR (Pineault & Joncas, 2000). HI observations from Kothes et al. (2001) place the system at 800 pc.
@@
X-ray observations show similar structures coincident with the head and tail regions. Fujita et al. (2021) found that the X-ray emission from the entire SNR is from nonthermal synchrotron radiation. However, the X-ray brightness in the head region contrasts with the gamma ray observations, which instead see a brightness in the tail region.
@@
Ground-based gamma ray particle detector arrays have so far only been able to resolve one emission spot in the Boomerang region. Milagro, VERITAS, Tibet, and LHAASO reported extended emission coincident with the tail of the SNR. Fermi-LAT (Fang et al., 2022) and MAGIC (Abe et al., 2022) reported gamma ray emission from the head and tail regions of the SNR separately, but no other gamma ray observations have been used to disentangle the sources thus far. Fang et al. (2022) separated the pulsar emission from the SNR and performed a detailed analysis of the tail region using Fermi-LAT data. They found that pion decay (PD) was the primary mechanism for producing the high-energy emission seen in the tail region. Abe et al. (2022) reported 0.16 extended gamma ray emission from the head and tail regions. Their multi-wavelength analysis revealed that the head region can have a leptonic or hadronic explanation, while the tail region is most likely of hadronic origin.
@@
The previous HAWC detection was analyzed using 1350 days of data and older reconstruction algorithms (Albert et al., 2020). Albert et al. (2020) found point-like emission consistent with the other gamma ray observations, which found extended emission in the tail region. The work presented in this paper uses more data and improved reconstruction algorithms (Albert et al., 2024), which further extends the previous HAWC measurement past 110 TeV with flux points. We also explore a molecular cloud template fit using data from the Planck 353 GHz dust opacity map and estimate the CR flux in the region using distance information from the Dame CO Survey (Dame et al., 2001). Section 2 describes the HAWC observatory and data used for this analysis. Section 3 describes the analysis procedure, as well as the results from the procedure. Gamma ray production mechanisms and their probable CR sources are described in section 4. Conclusions and further outlooks are presented in section 5.
@@ 2 HAWC Observatory and Data
The High-Altitude Water Cherenkov (HAWC) Observatory is located in Puebla, Mexico at an altitude of 4100 m. HAWC utilizes the water Cherenkov detection technique and consists of 300 water Cherenkov detectors (WCDs). Each WCD contains four photo-multiplier tubes (PMTs) and the entire array covers $\sim$22000 m2 (Abeysekara et al., 2023). For this analysis, we use 2565 days of data, which is about 1000 days more data than the previous publication Albert et al. (2020). A neural net (NN) algorithm is used to reconstruct the energy of the primary gamma ray that initiated the extensive air shower. From here, the data is then binned in a 2D binning scheme based on the fraction of PMTs triggered and reconstructed energy (Abeysekara et al., 2019). In addition to having more data, the new data reconstruction process uses algorithms that better optimize angular resolution.
@@ 3.1 Systematic Source Search
This first part of this analysis utilizes the Multi-Mission Maximum Likelihood (threeML) framework (Vianello et al., 2015), along with the HAWC Accelerated Likelihood (HAL) plugin (Abeysekara et al., 2022) to work through a step-by-step multi-source search method inspired by the Fermi-LAT Extended Source Catalog (Ackermann et al., 2017; Albert et al., 2023). ThreeML uses a likelihood ratio as a test statistic (TS) to reach a parameter set that maximizes the likelihood of the hypothesized model over the background-only model. The value of the TS is defined as follows:
@@
where Lalt is the model hypothesis and Lnull is the background only hypothesis. In order to arrive at the most statistically preferred model in our multi-source search method, we will be looking at the difference in TS between each nested model that is tested. There are two main components to this analysis described below: point source search and extension testing.
@@
    Add one point source with a power-law (PL) spectrum to the location of the highest TS value in the significance map, leaving the position and spectral values of the newly added source free while the position of the previous sources remains fixed.
@@
    Check the $\Delta$TS between the previous model and the most recently tested model:
@@
      If the $\Delta$TS $>$ 16, keep the source in the model and return to step 1/
@@
      If the $\Delta$TS $<$ 16, do not add the additional source to the model and stop adding point sources. Go to extended source testing.
@@
    Replace the highest TS point source with a 2D extended Gaussian and leave all other sources as point sources. Keep all locations fixed.
@@
    Check the $\Delta$TS between the previous model and the most recently tested model:
@@
      If the $\Delta$TS $>$ 16, accept the extended source model and go to step 3.
@@
      If the $\Delta$TS $<$ 16, reject the extended source model and move to the next highest TS value source. Go to step 1.
@@
    Check the TS values of the point sources in the model:
@@
      If the TS values of all point sources are $>$ 16, refit the model and float all parameters. Go to step 1 if there are untested point sources remaining. End the study if there are no more point sources remaining for extension testing.
@@
      If the TS of any point source is $<$ 16, remove the source(s) and refit the new model while floating all parameters. Go to step 1 if there are untested point sources remaining. End the study if there are no more point sources remaining for extension testing.
@@
After completing this procedure in the Boomerang region, the best-fitting model is found to be a single extended source with an LP spectrum. Table 1 lists the best-fit position and spectral parameters with their statistical and systematic uncertainties. The HAWC significance maps using the new reconstruction algorithms and best-fit values are shown in figure 2. The systematic uncertainties are determined using various detector response files that test possible effects on the data resulting from differences in the modeling of the instrument. These differences are calculated following Abeysekara et al. (2019).
[…]
@@ 3.2.1 Molecular Cloud Template Fitting
Since the Boomerang region is believed to be a PeVatron, the main production mechanism for the UHE gamma ray emission in the region is thought to be produced through PD. The head of the SNR houses the PSR and its PWN, which would mean that the protons are most likely being accelerated in the tail region where there is supernova (SN) ejecta and the shock front, which is coincident with a nearby molecular cloud. The molecular cloud is an ideal target for the accelerated protons to interact with and undergo PD. This means that UHE gamma ray morphology should take the shape of the molecular clouds that it is being produced in. This analysis explores fitting HAWC data above 56 TeV using a molecular cloud template from the Planck Collaboration (Ade et al., 2011) to explain a possible hadronic mechanism.
@@
Again, we utilize the threeML/HAL framework to do a maximum likelihood fitting of the molecular cloud spatial template. A similar procedure to that of Albert et al. (2021) is used to create the spatial template that is fed to threeML for the template fitting:
@@
    Calculate column density– this calculation is taken from Albert et al. (2021):
@@
    Crop data to the region of interest.
@@
    Calculate the mass of the region.
@@
    Normalize data to 1/sr– required for a 3ML 2D template
@@ 3.2.2 Simple Leptonic Model Fitting
We also explore the possibility of a leptonic mechanism producing the UHE HAWC data. This would mean that electrons are being accelerated by the SN ejecta or the PWN winds. For this scenario, we model the region with a point source morphology and a power-law spectrum, which has a fixed index of $-3.0$. Table 2 lists the best-fit values for the position and flux normalization after fitting the emission shown in figure 4.a with the simple model.
@@ 4.1 Cosmic Ray Energy Density from Protons
In section 3.2.1, we explore a hadronic-related fit to the HAWC UHE data with a molecular cloud template. We estimate the CR energy density using purely hadronic interactions and the measured gamma ray flux. The CR energy density can then be used to roughly determine the amount of energy that would be needed for the CR population in the region. We start by calculating the gamma ray energy flux using the best-fit gamma ray spectrum from our molecular cloud template fit:
@@
and $E_{i}$ = 56 TeV, $E_{f}$ = 316 TeV, and $k$, $E_{piv}$ and $\alpha$ are the same best-fit values from the fitting in section 3.2.1. Next, we use the integrated flux to calculate the luminosity:
@@
where d = 800 pc. Finally, we use the luminosity to calculate the CR proton energy density using the same approach as Abramowski et al. (2016):
@@
The SNR has a length of 14 pc and a width of 6 pc (Kothes et al., 2001). Since the tail is more elongated than the head, we assume that the tail is 2/3 of the length of the SNR. We also assume that the SNR is also capable of expanding 14 pc in all other directions as well, so we use a cube of 6 pc $\times$ 9.33 pc $\times$ 28 pc as the volume of the tail region as a higher estimate of the energy budget. Using these dimensions and the CR energy density, find that the energy budget for these CR protons that are $>$560 TeV is 7.62 $\times$ 1044 erg.
@@
4.2 Multi-Wavelength Modeling for Electrons
We use the results from section 3.2.2 to explore the possible leptonic nature of the UHE HAWC emission through multi-wavelength modeling. Since we assume that the hadronic nature of the UHE gamma ray emission is stemming from the shock fronts in the tail of the SNR, we assume the same scenario for the leptonic emission we model here. Therefore, we use other gamma ray observations that also model the tail region for additional information during our modeling. We use flux points from HAWC, MAGIC’s tail source (Abe et al., 2022), and Fermi-LAT (Fang et al., 2022) to model the gamma ray production using Naima (Zabalza, 2015). The X-ray observation is from the Suzaku telescope (Ge et al., 2021) and its flux is calculated using the ”East” region and a total solid angle of 17.82 arcmin2. The radio observations come from the DRAO at 408 MHz and 1240 MHz (Kothes et al., 2001), the Sino-German $\lambda$6 cm polarization survey (Gao et al., 2011), and the Effelsberg $\lambda$11 and $\lambda$21 cm surveys (Gao et al., 2011). Since the radio surveys for the Sino-German and Effelsberg surveys are reported for the entire SNR, their fluxes are scaled using the DRAO 408 MHz tail-to-whole flux ratio.
The multi-wavelength modeling was done using the Naima framework (Zabalza, 2015). All parent particle spectral models use a power-law with an exponential cutoff,
$f(E)=A(E/E_{0})^{-\alpha}\exp(-(E/E_{\mathrm{cutoff}})^{\beta})$ (9)
The gamma ray observations are assumed to be produced through Inverse Compton scattering (IC). For IC, the target photon fields considered are the cosmic microwave background (CMB) and near-infrared (NIR) (same as that used in Amenomori et al. (2021)). The radio and X-ray observations are assumed to be produced through nonthermal synchrotron radiation. This modeling also assumes a single population of electrons that are responsible for the radio, X-ray, and gamma ray observations. The best-fit values from the leptonic fit can be found in table 3 and figure 5 shows the multi-wavelength model.
Figure 5: Top: Multi-wavelength SED. The data shown includes the DRAO flux points in red squares (Pineault & Joncas, 2000), the Sino-German and Effelsberg measurements in dark green circles (Gao et al., 2011), the Suzaku spectrum in yellow (Ge et al., 2021), the Fermi-LAT flux points in black dots (Fang et al., 2022), the MAGIC flux points in green diamonds (Abe et al., 2022), and the HAWC flux points from the VHE work in blue triangles. Synchrotron is shown in dashed light grey, IC is shown in dotted dark grey, and black shows the total SED from synchrotron and IC combined. Bottom: Offset between the data points used and the model determined with Naima. The Fermi-LAT upper limits are removed to better see the deviation about zero for the other data points.
@@ 4.2 Multi-Wavelength Modeling for Electrons
The leptonic modeling in this section assumes electrons are producing the UHE HAWC data which, after being accelerated in the shock fronts of the SNR ejecta or the PWN’s wind, interact with the nearby photon fields and undergo IC scattering. The synchrotron and IC emission can explain the emission being seen in this region well (figure 5). Using this multi-wavelength model, we get a total electron energy budget of 1.28× $\leavevmode\nobreak\ \times\leavevmode\nobreak\ 10^{45}$ erg for electrons $>$1 TeV, which is below the energy budget estimation of the SNR.
While the PWN would be an ideal electron accelerator, it is offset from the location of the UHE HAWC emission and located in the head of the SNR, which we are not assuming the UHE gamma ray emission comes from. This would mean that electrons are accelerated in the shock fronts of the SN ejecta and produce gamma rays through IC. Since we are only using energy as our differentiator for the head and tail regions of the SNR, it is also possible that electrons accelerated from the SN ejecta are producing most of the gamma ray emission, but some ”leakage” from the PWN is also being accounted for in the UHE gamma ray emission.

### v2

@@ Abstract
Aims. We aim to characterize the very-high-energy (VHE, 0.1–100 TeV) gamma ray emission from SNR G106.3+2.7 by determining the morphology and spectral energy distribution of the region. This is accomplished using 2565 days of data and improved reconstruction algorithms from the High Altitude Water Cherenkov (HAWC) Observatory. We also explore possible gamma ray production mechanisms for different energy ranges.
@@ 1 Introduction
SNR G106.3+2.7 is a PeVatron candidate that has been under study for many years now. It was first detected in 1990 by Joncas & Higgs (1990) at 408 MHz with the Dominion Radio Astrophysical Observatory (DRAO). The head of the SNR is home to the pulsar J2229+6114 and its boomerang-like pulsar wind nebula (PWN) (Kothes et al., 2001), known as the Boomerang region. An extended shell then reaches into the tail region of the SNR (Pineault & Joncas, 2000). There have been many proposed distances for this region. Pineault & Joncas (2000) adopted a distance of 12 kpc for the SNR based on a flat rotation curve and Halpern et al. (2001) found the distance to be 3 kpc using X-ray absorption column density. HI observations from Kothes et al. (2001) place the system at 800 PC.
@@
X-ray observations show similar structures coincident with the head and tail regions. Fujita et al. (2021) finds that the X-ray emission from the entire SNR is from nonthermal synchrotron radiation. However, the X-ray brightness in the head region contrasts with the gamma ray observations, which instead see a brightness in the tail region.
@@
Ground-based gamma ray particle detector arrays have so far only been able to resolve one emission spot in the Boomerang region. Milagro, VERITAS, Tibet, and LHAASO report extended emission coincident with the tail of the SNR. Fermi-LAT (Fang et al., 2022) and MAGIC (Abe et al., 2022) report gamma ray emission from the head and tail regions of the SNR separately, but no other gamma ray observations have been used to disentangle the sources thus far. Fang et al. (2022) separates the pulsar emission from the SNR and performs a detailed analysis of the tail region using Fermi-LAT data. They find that pion decay (PD) is the primary mechanism for producing the high-energy emission seen in the tail region.
Abe et al. (2022) reports 0.16 extended gamma ray emission from the head and tail regions. They also show that the emission shifted in the SNR depending on the energy threshold. Their highest energy emission (6.0-30 TeV) is consistent with the tail of the SNR. Their multi-wavelength analysis reveals that the head region can have a leptonic or hadronic explanation, while the tail region is most likely of hadronic origin.
@@
The previous HAWC detection analyzes 1350 days of data using older reconstruction algorithms (Albert et al., 2020). Albert et al. (2020) finds point-like emission consistent with the other gamma ray observations, which see extended emission in the tail region. The work presented in this paper uses more data and improved reconstruction algorithms (Albert et al., 2024), which further extends the previous HAWC measurement past 110 TeV with flux points. We also explore a molecular cloud template fit using data from the Planck 353 GHz dust opacity map and estimate the CR flux in the region using distance information from the Dame CO Survey (Dame et al., 2001). Section 2 describes the HAWC observatory and data used for this analysis. Section 3 describes the analysis procedure, as well as the results from the procedure. Gamma ray production mechanisms and their probable CR sources are described in section 4. Conclusions and further outlooks are presented in section 5.
@@ 2 HAWC Observatory and data
The High-Altitude Water Cherenkov (HAWC) Observatory is located in Puebla, Mexico at an altitude of 4100 m. HAWC utilizes the water Cherenkov detection technique and consists of 300 water Cherenkov detectors (WCDs). Each WCD contains four photo-multiplier tubes (PMTs) and the entire array covers $\sim$22000 m2 (Abeysekara et al., 2023). For this analysis, we used 2565 days of data, which is about 1000 days more data than the previous publication Albert et al. (2020). A neural net (NN) algorithm was used to reconstruct the energy of the primary gamma ray that initiated the extensive air shower. From here, the data was binned in a 2D binning scheme based on the fraction of PMTs triggered and reconstructed energy (Abeysekara et al., 2019). In addition to having more data, the new data reconstruction process used algorithms that better optimize angular resolution.
@@ 3.1 Systematic source search
[…]
This first part of this analysis utilized the Multi-Mission Maximum Likelihood (threeML) framework (Vianello et al., 2015), along with the HAWC Accelerated Likelihood (HAL) plugin (Abeysekara et al., 2022) to work through a step-by-step multi-source search method inspired by the Fermi-LAT Extended Source Catalog (Ackermann et al., 2017; Albert et al., 2023). ThreeML uses a likelihood ratio as a test statistic (TS) to reach a parameter set that maximizes the likelihood of the hypothesized model over the background-only model. The value of the TS is defined as follows:
@@
where Lalt is the model hypothesis and Lnull is the background only hypothesis. In order to arrive at the most statistically preferred model in our multi-source search method, we looked at the difference in TS between each nested model that is tested. There are two main components to this analysis described below: point source search and extension testing.
@@
    The first step of our multi-source search is to add one point source with a power-law (PL) spectrum to the location of the highest TS value in the significance map, leaving the position and spectral values of the newly added source free while the position of the previous sources remains fixed.
@@
    After fitting the model in step 1, we check the $\Delta$TS between the previous model and the most recently tested model:
@@
      If the $\Delta$TS $>$ 25, we keep the source in the model and return to step 1.
@@
      If the $\Delta$TS $<$ 25, we do not add the additional source to the model and stop adding point sources. We then start our extended source testing.
@@
    We start by replacing the highest TS point source with a 2D extended Gaussian and leave all other sources as point sources. We keep all source locations fixed.
@@
    After fitting the model in step 1, we check the $\Delta$TS between the previous model and the most recently tested model:
@@
      If the $\Delta$TS $>$ 16, we accept the extended source model and go to step 3.
@@
      If the $\Delta$TS $<$ 16, we reject the extended source model and move to the next highest TS value source. We then start back at step 1.
@@
    Next, we check the TS values of the point sources in the model:
@@
      If the TS values of all point sources are $>$ 25, we refit the model and float all parameters. If untested point sources remain, we go back to step 1. We end the study if there are no more point sources remaining for extension testing.
@@
      If the TS of any point source is $<$ 25, we remove the source(s) and refit the new model while floating all parameters. If untested point sources remain, we go back to step 1. We end the study if there are no more point sources remaining for extension testing.
@@
After completing this procedure in the Boomerang region, the best-fitting model was found to be a single extended source with an LP spectrum. Table 1 lists the best-fit position and spectral parameters with their statistical and systematic uncertainties. The HAWC significance maps using the new reconstruction algorithms and best-fit values are shown in figure 2. The systematic uncertainties were determined using various detector response files that test possible effects on the data resulting from differences in the modeling of the instrument. These differences are calculated following Abeysekara et al. (2019).
@@ 3.2 HAWC data $>$56 TeV
We will now explore the possible production mechanisms that would contribute to the PeVatron’s UHE emission in this region. Most PeVatron theory assumes that these systems are hadronic, which is confirmed by correlation with molecular clouds in the area. However, recent studies suggest that the leptonic scenario is more than capable of producing UHE emission as well (Vannoni et al., 2009; Breuhaus et al., 2021). Both scenarios will be considered for this region.
@@ 3.2.1 Molecular cloud template fitting
We will first look at the UHE gamma ray emission in the region using the pion decay production mechanism. The head of the SNR houses the pulsar and its PWN, which would mean that the protons are most likely being accelerated in the tail region where there is supernova (SN) ejecta and the shock front, which is coincident with a nearby molecular cloud. The molecular cloud is an ideal target for the accelerated protons to interact with and undergo PD. This means that UHE gamma ray morphology should take the shape of the molecular clouds that it is being produced in. This analysis explores fitting HAWC data above 56 TeV using a molecular cloud template from the Planck Collaboration (Ade et al., 2011) to explain a possible hadronic mechanism.
@@
Again, we utilized the threeML/HAL framework to do a maximum likelihood fitting of the molecular cloud spatial template. A similar procedure to that of Albert et al. (2021) was used to create the spatial template that is fed to threeML for the template fitting:
@@
    We start by calculating the column density– this calculation is taken from Albert et al. (2021):
@@
    We then crop the data to the size of the region of interest.
@@
    We calculate the mass of the region using the following equation:
@@
    Finally, we normalize the data to 1/sr. This is required for a 3ML 2D template.
@@ 3.2.2 Simple leptonic model fitting
We also explore the possibility of a leptonic mechanism producing the UHE HAWC data. This would mean that electrons are being accelerated by the SN ejecta or the PWN winds. For this scenario, we modeled the region with a point source morphology and a power-law spectrum, which has a fixed index of $-3.0$. Table 2 lists the best-fit values for the position and flux normalization after fitting the emission shown in figure 4.a with the simple model.
@@ 4.1 Cosmic ray energy density from protons
In section 3.2.1, we explored a hadronic-related fit to the HAWC UHE data with a molecular cloud template. We estimated the CR energy density using purely hadronic interactions and the measured gamma ray flux. The CR energy density can then be used to roughly determine the amount of energy that would be needed for the CR population in the region. We started by calculating the gamma ray energy flux using the best-fit gamma ray spectrum from our molecular cloud template fit:
@@
and $E_{i}$ = 56 TeV, $E_{f}$ = 316 TeV, and $k$, $E_{piv}$ and $\alpha$ are the same best-fit values from the fitting in section 3.2.1. Next, we used the integrated flux to calculate the luminosity:
@@
where d = 800 pc. Finally, we used the luminosity to calculate the CR proton energy density using the same approach as Abramowski et al. (2016):
@@
The SNR has a length of 14 pc and a width of 6 pc (Kothes et al., 2001). Since the tail is more elongated than the head, we assumed that the tail is two-thirds of the length of the SNR. We also assumed that the SNR is also capable of expanding 14 pc in all other directions as well, so we use a cube of 6 pc $\times$ 9.33 pc $\times$ 28 pc as the volume of the tail region as a higher estimate of the energy budget. Using these dimensions and the CR energy density, we found that the energy budget for these CR protons that are $>$560 TeV is 7.62 $\times$ 1044 erg.
@@
4.2 Multi-wavelength modeling for electrons
We used the results from section 3.2.2 to explore the possible leptonic nature of the UHE HAWC emission through multi-wavelength modeling. Since we assumed that the hadronic nature of the UHE gamma ray emission stems from the shock fronts in the tail of the SNR, we assumed the same scenario for the leptonic emission we model here. Therefore, we used other gamma ray observations that also modeled the tail region for additional information during our modeling. We used flux points from HAWC, MAGIC’s tail source (Abe et al., 2022), and Fermi-LAT (Fang et al., 2022) to model the gamma ray production using Naima (Zabalza, 2015). The X-ray observation is from the Suzaku telescope (Ge et al., 2021) and its flux was calculated using the ”East” region and a total solid angle of 17.82 arcmin2. The radio observations come from the DRAO at 408 MHz and 1240 MHz (Kothes et al., 2001), the Sino-German $\lambda$6 cm polarization survey (Gao et al., 2011), and the Effelsberg $\lambda$11 and $\lambda$21 cm surveys (Gao et al., 2011). Since the radio surveys for the Sino-German and Effelsberg surveys are reported for the entire SNR, their fluxes were scaled using the DRAO 408 MHz tail-to-whole flux ratio.
The multi-wavelength modeling was done using the Naima framework (Zabalza, 2015). All parent particle spectral models used a power-law with an exponential cutoff,
$f(E)=A(E/E_{0})^{-\alpha}\exp(-(E/E_{\mathrm{cutoff}})^{\beta})$ (9)
The gamma ray observations are assumed to be produced through IC. For IC, the target photon fields considered are the cosmic microwave background (CMB) and near-infrared (NIR) (same as that used in Amenomori et al. (2021)). The radio and X-ray observations are assumed to be produced through nonthermal synchrotron radiation. This modeling also assumes a single population of electrons that are responsible for the radio, X-ray, and gamma ray observations. The best-fit values from the leptonic fit can be found in table 3 and figure 5 shows the multi-wavelength model.
Figure 5: Top: Multi-wavelength SED. The data shown includes the DRAO flux points in red squares (Pineault & Joncas, 2000), the Sino-German and Effelsberg measurements in dark green circles (Gao et al., 2011), the Suzaku spectrum in yellow (Ge et al., 2021), the Fermi-LAT flux points in black dots (Fang et al., 2022), the MAGIC flux points in green diamonds (Abe et al., 2022), and the HAWC flux points from the VHE work in blue triangles. Synchrotron is shown in dashed light grey, IC scattering is shown in dotted dark grey, and black shows the total SED from synchrotron and IC scattering combined. Bottom: Offset between the data points used and the model determined with Naima. The Fermi-LAT upper limits are removed to better see the deviation about zero for the other data points.
The leptonic modeling in this section assumed electrons are producing the UHE HAWC data which, after being accelerated in the shock fronts of the SNR ejecta or the PWN’s wind, interact with the nearby photon fields and undergo IC scattering. The synchrotron and IC emission can explain the emission being seen in this region well (figure 5). Using this multi-wavelength model, we got a total electron energy budget of 1.28× $\leavevmode\nobreak\ \times\leavevmode\nobreak\ 10^{45}$ erg for electrons $>$1 TeV, which is below the energy budget estimation of the SNR.
Breuhaus et al. (2021) suggests that electron cooling dominated by Inverse Compton (IC) losses could produce UHE emission, resulting in a PeVatron. One requirement for this system is a pulsar with a spin-down energy of $L_{SD}>10^{36}$ erg/s. While the PWN in this region fits that requirement ($L_{SD}\approx 2\times 10^{37}$ erg/s) and would be an ideal electron accelerator, it is offset from the location of the UHE HAWC emission and located in the head of the SNR, which we are not assuming the UHE gamma ray emission comes from. This would mean that electrons are accelerated in the shock fronts of the SN ejecta and produce gamma rays through IC.
As mentioned in section 1, MAGIC also saw a shift in emission towards the tail of the SNR as energy thresholds increased (Abe et al., 2022). Even though HAWC’s angular resolution is not as small as MAGIC’s, we do see our UHE emission trend towards the tail of the SNR (figure 4.b). Since we are only using energy as our differentiator for the head and tail regions of the SNR, it is also possible that electrons accelerated from the SN ejecta are producing most of the gamma ray emission, but some ”leakage” from the PWN is also being accounted for in the UHE gamma ray emission.